OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!! An expression is shown below: f(x) = -16x2 + 24x + 16 What are the x-intercepts of the graph of f(x)? Show your work.

OpenStudy (anonymous):

@SolomonZelman @jim_thompson5910 @mathmale @satellite73

OpenStudy (anonymous):

tranlate as solve for \(x\) :\[-16x^2 + 24x + 16 =0\]

OpenStudy (anonymous):

i would divide by \(-8\) and start with \[2x^2-3x-2=0\] if you are lucky it factors, elsewise use the quadratic formula

OpenStudy (anonymous):

Ok, so I got (x-2)(2x+1) = 0

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

set each factor equal to zero and solve for \(x\) for example \[x-2=0\iff x=2\]

OpenStudy (anonymous):

ok, so is the answer 2 and -.5 ?

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks for your help! :)

OpenStudy (anonymous):

yw